Can someone please help:
Write the equation of the circle with the given center and length of a radius. Express the final equation in the form X^2+Y^2+Dx+Ey+F=0
cCenter at (3,5) and r=2
Thanks
Charmaine

Answer by Nate(3495)

the formula is:
(x - h)^2 + (y - k)^2 = r^2 ............. center(h,k) and r is radius
(x - 3)^2 + (y - 5)^2 = 4
x^2 - 6x + 9 + y^2 - 10y +25 = 4
x^2 + y^2 - 6x - 10y + 30 = 0